# Benchmarking and Improving Bird’s Eye View Perception Robustness in Autonomous Driving

### v1

@@ 6.4 Backbone
- The Swin Transformer is more vulnerable towards the lighting changings; VoVNet-V2 is more robust against Snow while ResNet shows better robustness across a wide range of corruptions. Although ResNet [120] and VoVNet [63] exhibit close standard performance, ResNet-based detectors exhibit consistently superior robustness across a wide range of corruptions, as illustrated in Figure 8. Conversely, the VoVNet backbone consistently exhibits better robustness under Snow corruptions. Moreover, Swin Transformer [45] based BEVDet [3] demonstrates significant vulnerability towards changes in lighting conditions (e.g., Bright and Dark). A clear comparison can be found in Figure 9. Inspired by [119], we compute the Gramian matrix in the feature space in the features extracted by the backbone model, under benign inputs and corruption inputs. Based on that, we compute the relative error between the “clean” and corruption inputs Gramian matrix. The VoVNet has a relative error of 4.17 and 5.43 towards Snow and Motion Blur, respectively. While ResNet has errors of 11.47 and 2.10. The results align with our end-to-end experiments above, where ResNet is more robust under Motion Blur while VoVNet is more robust towards Snow. The observed robustness differences between ResNet and VoVNet to Snow and Motion Blur stem from their architectural designs. ResNet’s deep hierarchical features and skip connections make it more sensitive to high-frequency noise like Snow, resulting in higher errors. Conversely, VoVNet’s dense feature aggregation and efficient feature use enhance robustness to Snow by mitigating local noise patterns. However, VoVNet shows higher errors under Motion Blur. ResNet’s hierarchical feature representation might be better suited to handling the spatial inconsistencies caused by simulated Motion Blur. These findings suggest VoVNet excels in spatial noise resilience, while ResNet handles spatial distortions better. Furthermore, the finding can provide insights for feature work to design new loss functions towards aligning the Gramian matrix of features extracted by backbone under different corruptions.
@@ 7 Conclusion
Potential Limitation. Despite the distinct data corruptions we introduce in our benchmarks, they cannot cover all the out-of-distribution scenarios in real-world applications due to unpredictable complexity. Additionally, we mainly analyzed coarse-grained designs among models (e.g., depth estimation) since it is considerably non-trivial to identify the trade-off between fine-grained network designs.
Acknowledgments. This work is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221- 0012), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s). This work is also under the programme DesCartes and is supported by the National Research Foundation, Prime Minister’s Office, Singapore, under its Campus for Research Excellence and Technological Enterprise (CREATE) programme.

### v2

@@ 6.4 Backbone
- The Swin Transformer is more vulnerable towards the lighting changings; VoVNet-V2 is more robust against Snow while ResNet shows better robustness across a wide range of corruptions. Although ResNet [120] and VoVNet [63] exhibit close standard performance, ResNet-based detectors exhibit consistently superior robustness across a wide range of corruptions, as illustrated in Figure 8. Conversely, the VoVNet backbone consistently exhibits better robustness under Snow corruptions. Moreover, Swin Transformer [45] based BEVDet [3] demonstrates significant vulnerability towards changes in lighting conditions (e.g., Bright and Dark). A clear comparison can be found in Figure 9. Inspired by [119], we compute the Gramian matrix in the feature space in the features extracted by the backbone model, under clean inputs and corruption inputs. Based on that, we compute the relative error between the “clean” and corruption inputs Gramian matrix. The VoVNet has a relative error of 4.17 and 5.43 towards Snow and Motion Blur, respectively. While ResNet has errors of 11.47 and 2.10. The results align with our end-to-end experiments above, where ResNet is more robust under Motion Blur while VoVNet is more robust towards Snow. The observed robustness differences between ResNet and VoVNet to Snow and Motion Blur stem from their architectural designs. ResNet’s deep hierarchical features and skip connections make it more sensitive to high-frequency noise like Snow, resulting in higher errors. Conversely, VoVNet’s dense feature aggregation and efficient feature use enhance robustness to Snow by mitigating local noise patterns. However, VoVNet shows higher errors under Motion Blur. ResNet’s hierarchical feature representation might be better suited to handling the spatial inconsistencies caused by simulated Motion Blur. These findings suggest VoVNet excels in spatial noise resilience, while ResNet handles spatial distortions better. Furthermore, the finding can provide insights for feature work to design new loss functions towards aligning the Gramian matrix of features extracted by backbone under different corruptions.
@@ 7 Conclusion
Potential Limitation. Despite the distinct data corruptions we introduce in our benchmarks, they cannot cover all the out-of-distribution scenarios in real-world applications due to unpredictable complexity. Additionally, we mainly analyzed coarse-grained designs among models (e.g., depth estimation) since it is considerably non-trivial to identify the trade-off between fine-grained network designs. Furthermore, we do not include joint corruption on camera and LiDAR, which is essential for some real-world scenarios (e.g., the Snow and Fog weather cases as shown in Table V).